# DeepPAAC: A New Deep Galerkin Method for Principal-Agent Problems

### v1

@@ 1 Introduction
Despite the large volume of work, strongly focused on explicit solutions, few articles investigate numerical algorithms for PA problems. Recently, several papers utilized machine learning methods to do so [baldacci2019market, campbell2021deep]. Baldacci et al. [baldacci2019market] adopt an actor–critic reinforcement learning approach that learns the control and contract explicitly via temporal-difference updates and policy gradients; Campbell et al. [campbell2021deep] discretize an FBSDE and train a deep–BSDE solver by minimizing the terminal mismatch for the inner problem. Compared to these works, we solve the HJB equation by directly minimizing the PDE residual at space–time samples under boundary and terminal conditions, thus requiring neither time discretization nor FBSDE simulations.
@@
We utilize the ideas in [al2019applications] and develop a new neural network architecture inspired by the swish activation function proposed in [ramachandran2017searching] and a direct encoding of the terminal constraint proposed in [sukumar2022exact]. Our DeepPAAC scheme solves the Principal’s HJB equation by minimizing the loss function of the differential operator, first order condition of the control, as well as a penalty function of the control (for constrained Principal/Agent controls) with a mesh free approach. Numerical results in our paper show that the algorithm is stable, accurate and fast, outperforming the traditional DGM neural network.
@@ 2 The Model
We follow the formulation in [Sung2025] to describe a general model for continuous time Principal-Agent (PA) problems.
@@
where $U^{A}$ and $\Phi^{A}$ are both increasing, concave utility functions in $\alpha,\beta,\xi$ respectively. $\mathbb{E}^{a}$ means the expectation is under the probability $\mathbb{P}^{a}$ mentioned above. Given $\{\alpha_{t},\beta_{t}\}_{\{0\leq t\leq T\}}$ and $\xi$, the Agent maximizes his utility by choosing optimal $\{a_{t}\}_{\{0\leq t\leq T\}}$ and $\{c_{t}\}_{\{0\leq t\leq T\}}$:
@@
We denote by $a(s,\alpha,\beta,\xi)$ and $c(s,\alpha,\beta,\xi)$ the optimal controls achieving the supremum above, viewed as best response functions to the Principal’s choices of $(\alpha,\beta,\xi)$. By convention, if the Agent has multiple optimal controls $a(s,\alpha,\beta,\xi)$ and $c(s,\alpha,\beta,\xi)$, it is assumed that the Agent always chooses the one that is best for the Principal.
@@
$J^{P}(t,\alpha,\beta,\xi,D)=\mathbb{E}^{a(s,\alpha,\beta,\xi)}\left[\int_{t}^{T}U^{P}(s,\alpha_{s},\beta_{s},a(s,\alpha,\beta,\xi),D_{s})ds+\Phi^{P}(T,\xi,m^{P}_{T})\right],$ (2.7)
where $U^{P}$ and $\Phi^{P}$ are both increasing, concave utility functions in $D,m^{P}$ respectively. Observe that the expectation in (2.7) is based on the feedback probability measure $\mathbb{P}^{a(s,\alpha,\beta,\xi)}$. The Principal’s problem at time $t$ is
@@ 2.1 HJB Equations and Critic Actor Algorithm
For the sake of simplification (by setting $\tilde{m}^{A}_{t}=e^{-r_{A}t}m_{t}^{A}$ and $\tilde{m}^{P}_{t}=e^{-r_{p}t}m_{t}^{P}$), we may assume zero interest rates $r_{A}=r_{p}=0$ and to capture the most important feature of Principal-Agent problem, we choose $b(t,a,x)=a,\sigma(t,a,x)=1$. The PA problem constitutes a stochastic control problem; hence, its corresponding Principal’s value function admits characterization via an associated Hamilton-Jacobi-Bellman (HJB) equation. As is standard in the literature (e.g. [sannikov2008continuous]), we first solve the Agent’s problem (2.6) by using the comparison principle of BSDEs and stochastic maximum principle. When solving the Agent’s problem, the Agent’s optimal remaining utility process is characterized as a state process, which we denote as $\{W_{t}\}_{\{0\leq t\leq T\}}$. An auxiliary process $\{Y_{t}\}_{\{0\leq t\leq T\}}$ will also arise when deriving the Agent’s optimal consumption process, . These two processes, $\{W_{t}\}_{\{0\leq t\leq T\}}$, $\{Y_{t}\}_{\{0\leq t\leq T\}}$, together with (2) and (2.4), are the four state processes needed to solve the Principal’s problem.
These states follow the following dynamics for $t\leq s\leq T$:
$\begin{cases}W_{s}&=w-\int_{t}^{s}U^{A}(r,\alpha_{r},\beta_{r},I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r}),I_{c}(\alpha_{r},\beta_{r},Y_{r},Z_{r}))dr+\int_{t}^{s}Z_{r}dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\cr Y_{s}&=y+\int_{t}^{s}\hat{Z}_{r}dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\cr m^{A}_{s}&=\int_{t}^{s}[\alpha_{r}-I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})+\beta_{r}I_{c}(\alpha_{r},\beta_{r},Y_{r},Z_{s})]ds+\int_{t}^{s}\beta_{r}dB^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r}^{A})}_{r},\cr m^{P}_{s}&=\int_{t}^{s}((1-\beta_{r})I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})-D_{r}-\alpha_{r})dr+\int_{t}^{s}(1-\beta_{r})dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\end{cases}$ (2.9)
@@
$V(t,w,y,0,0)=\sup_{\alpha,\beta,D,Z,\hat{Z}}\mathbb{E}^{I_{a}(\alpha,\beta,Y,Z)}\left[\int_{t}^{T}U^{P}(s,\alpha_{s},\beta_{s},I_{a}(\alpha_{s},\beta_{s},Y_{s},Z_{s}),D_{s})ds+\Phi^{P}(T,\xi,m^{P}_{T})\right],$
@@ 2.2 Policy Iteration
In the past, such problems were viewed as very challenging, since they require multiple levels of approximation—locally to solve the supremum and then globally to solve for $V(t,x)$. Below we develop a deep-learning based setup to address this aspect of PA HJB. The key idea that we rely on is to construct (overparametrized) functional approximator that replaces $V(t,x)$ with a surrogate $\hat{V}(t,x)$ that is trained through gradient descent. The approximation is global in $(t,x)$ (e.g. there is no finite-differencing) and moreover carries through to the derivatives, which are evaluated using algorithmic differentiation. In turn, functional approximation allows to do the same for the optimal control $u(\cdot,\cdot)$, which is similarly approximated via a deep neural network.
The above construction is conceptually dimension-agnostic, and moreover thanks to being overparametrized can be trained exhaustively, breaking down the overall solver into a loop consisting of taking stochastic gradient descents over training batches that are refreshed across epochs. Consequently, the method has modest memory requirements and does not rely on any temporal or spatial discretization.
@@
with unknown value function $V$ and unknown optimal feedback control ${\bf u}\equiv u(t,{\bf x})$, with ${\cal L}^{\bf u}V(t,{\bf x})\equiv{\cal L}^{u(t,{\bf x})}V(t,{\bf x})$. Here, both $\bf x$} and $\bf u$} are vectors in ${\cal X}\subseteq\mathbb{R}^{d}$ and ${\cal U}\subseteq\mathbb{R}^{a}$ respectively.
@@ 3 Implementation
3.1 Loss Function
Due to the presence of multiple convergence criteria, determining a loss function of our PAAC algorithm is nontrivial.
We define
$L_{ctrl}(\theta_{n}^{V};t_{n}^{m},{\bf x}_{n}^{m})=({\nabla{\cal L}\over\nabla u})^{u(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u})}v(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{V})$
to verify whether step 6 in our PAAC algorithm is converging (note that $\bf u$} is vector, hence $L_{ctrl}(\theta_{n}^{V};t_{n}^{m},{\bf x}_{n}^{m})$ is also a vector). To define a stopping criterion, we construct a validation set $\{\bar{t}^{m},\bar{{\bf x}}^{m}\},m=1,\ldots,M_{V}$. Unlike the training batches that vary across steps, the validation set is fixed throughout, to allow an objective tracking of surrogate quality and algorithm convergence and to mitigate over-fitting to a particular training epoch. For the stopping criterion, we seek to control the $L_{\infty}$ norm of the vectors $\bar{L}_{int}$ and $\bar{L}_{ctrl}$ which are counterparts of $L_{int},{L}_{ctrl}$ over the validation set. Thus, we continue the PAAC iterations until we ensure that the maximum pointwise errors of $v_{n},u_{n}$ in the interior are small:
$\|\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})\|_{\infty}\leq\mathrm{Tol_{int}},\qquad\|\bar{L}_{ctrl}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})\|_{\infty}\leq\mathrm{Tol_{ctrl}},$
@@ 3.2 Neural Network Architecture
We consider two different architectures for the neural networks. The first one is the DGM (Deep Galerkin Method) architecture used by Sirignano and Spiliopoulos [sirignano2018dgm]. The second one is a standard fully-connected feed-forward neural network.
@@
To evaluate the differential operator ${\cal L}^{u(\cdot;\theta^{u}_{n})}$ that defines the loss $L_{int}$ we compute numerical gradients of the current value function approximation $v(\cdot;\theta^{V}_{n})$ using backpropagation. To this end we rely on the native backpropagation (aka algorithmic differentiation) support in modern ML libraries, such as the function tf.gradients in Python’s Tensorflow library. This capability together with the underlying universal approximation property of deep NN’s are the motivating factors for using NN approximators, as it allows to replace “term-for-term” the HJB equation with a corresponding neural network approximation.
@@ 3.3 Training Designs
As a third option, to enhance convergence, we consider adaptive sampling, where new training points are picked based on the current loss $L_{int}(\cdot,\cdot)$, targeting regions where the current loss is highest. This achieves two purposes. First, the use of additional training locations helps to achieve global minimization of the loss metric. Recall that $L_{n}$ is supposed to represent the integrated error over the entire domain; the approximation using a fixed set of $(t^{m},{\bf x}^{m})$ runs the danger of overfitting at those specific sites. Second, by preferentially picking $(t^{m},{\bf x}^{m})$ we can speed-up convergence by focusing on regions where the present fit is less good.
@@ 3.4 A toy example: The Holmström–Milgrom Model
with exponential utility $\Phi^{A}(x)=-\exp(-\gamma_{A}x)$ for some positive $\gamma_{A}>0$. We denote by $a(s,\xi)$ the best response in the Agent’s problem. Solving the latter (see proof of Proposition 3.1), we have the terminal payment expressed as
@@
The first order condition is:
${\cal F}^{Z}:=V_{x}^{P}[1-(1+\gamma_{A})Z]+V_{xx}^{P}(Z-1).$ (3.23)
@@
Figure 2: Estimated optimal control $\hat{a}(t,x)$ obtained from the DeepPAAC algorithm for different exponential mixture weights $\lambda\in(0,1)$.
@@
Figure 3: Left: $\hat{a}(t,x;0)$ for $\lambda=0$. Right: $\hat{a}(t,x;0.5)$ for $\lambda=0.5$
For neural network surrogates of the value function $V(\cdot,\cdot)$ and the control $a(\cdot,\cdot)$, we use our DeepPAAC architecture described in Section 3.2. In the illustration below we take $L=3$ layers, with 32 neurons per layer. The networks $\theta_{0}^{V},\theta_{0}^{u}$ are initialized using Glorot uniform initialization for the weights and zeros initialization for the biases. The stopping rule is $\mathrm{Tol_{int}}=\mathrm{Tol_{ctrl}}=10^{-3}$. For training we use batches of size $M=2000$, polynomial learning rate decay from $10^{-3}$ to $10^{-4}$ with power $0.8$. In the given example , we need 6750/6600 steps to satisfy the stopping criterion.
In Figure 2, we plot the numerically obtained optimal control $\hat{a}(t,x;\lambda)$ at a few locations $(t,x)=({1\over 4},{1\over 4}),({1\over 2},{1\over 2}),({3\over 4},{3\over 4})$ as we vary the mixture weight $\lambda$. The parameters in Figure 2 are $\gamma_{A}=0.5$, $\gamma_{P}=1$ and $\tilde{\gamma}_{P}=2$. We observe that $\hat{a}(t,x;\lambda)>a^{Lin}(t,x;\lambda)$ while $\hat{a}(t,x;0)$ and $\hat{a}(t,x;1)$ match the analytical exact solution. The left panel of Figure 3 shows $\hat{a}(t,x;0)$ as a function of $t,x$. As expected, the NN surrogate correctly learns that $\hat{a}(t,x;0)$ is (almost) a constant. The right panel of Figure 3 repeats this for $\hat{a}(t,x;0.5)$. We observe that when $\lambda=0.5$, $\hat{a}(t,x;\lambda)$ increases in $t$ and decreases in $x$.

### v2

@@ 1 Introduction
Despite the large volume of work, strongly focused on explicit solutions, few articles investigate numerical algorithms for PA problems. Recently, several papers utilized machine learning methods to do so [baldacci2019market, campbell2021deep, dayanikli2024machine]. Baldacci et al. [baldacci2019market] adopt an actor–critic reinforcement learning approach that learns the control and contract explicitly via temporal-difference updates and policy gradients; Campbell et al. [campbell2021deep] discretize an FBSDE and train a deep–BSDE solver by minimizing the terminal mismatch for the inner problem. Dayanıklı and Lauriere [dayanikli2024machine] use the method of Lagrange multipliers to rewrite the cost function and discretize the SDE under the framework of mean field game. Compared to these works, we solve the HJB equation by directly minimizing the PDE residual at space–time samples under boundary and terminal conditions, thus requiring neither time discretization nor FBSDE simulations.
@@
We utilize the ideas in [al2019applications, al2022extensions] and develop a new neural network architecture inspired by the swish activation function proposed in [ramachandran2017searching] and a direct encoding of the terminal constraint proposed in [sukumar2022exact]. Our DeepPAAC scheme solves the Principal’s HJB equation by minimizing the loss function of the differential operator, first order condition of the control, as well as a penalty function of the control (for constrained Principal/Agent controls) with a mesh free approach. Numerical results in our paper show that the algorithm is stable, accurate and fast, outperforming the traditional DGM neural network.
@@ 2 The Model
We follow the formulation in [Sung2025] to describe a general model for continuous time Principal-Agent (PA) problems  [^1] Technical conditions are added in [Sung2025] to guarantee the well-posedness of the problem..
@@
where $U^{A}$ and $\Phi^{A}$ are both increasing, concave and integrable utility functions in $\alpha,\beta,\xi$ respectively. $\mathbb{E}^{a}$ means the expectation is under the probability $\mathbb{P}^{a}$ mentioned above. Given $\{\alpha_{t},\beta_{t}\}_{\{0\leq t\leq T\}}$ and $\xi$, the Agent maximizes his utility by choosing optimal $\{a_{t}\}_{\{0\leq t\leq T\}}$ and $\{c_{t}\}_{\{0\leq t\leq T\}}$:
@@
We denote by $a^{*}(s,\alpha,\beta,\xi)$ and $c^{*}(s,\alpha,\beta,\xi)$ the optimal controls achieving the supremum above, viewed as best response functions to the Principal’s choices of $(\alpha,\beta,\xi)$. By convention, if the Agent has multiple optima $a(s,\alpha,\beta,\xi)$ and $c(s,\alpha,\beta,\xi)$, it is assumed that the Agent always chooses the one that is best for the Principal.
@@
$J^{P}(t,\alpha,\beta,\xi,D)=\mathbb{E}^{a^{*}(s,\alpha,\beta,\xi)}\left[\int_{t}^{T}U^{P}(s,\alpha_{s},\beta_{s},a(s,\alpha,\beta,\xi),D_{s})ds+\Phi^{P}(T,\xi,m^{P}_{T})\right],$ (2.7)
where $U^{P}$ and $\Phi^{P}$ are both increasing, concave utility functions in $D,m^{P}$ respectively. Observe that the expectation in (2.7) is based on the feedback probability measure $\mathbb{P}^{a^{*}(s,\alpha,\beta,\xi)}$. The Principal’s problem at time $t$ is
@@ 2.1 HJB Equations and Critic Actor Algorithm
For the sake of simplification (by setting $\tilde{m}^{A}_{t}=e^{-r_{A}t}m_{t}^{A}$ and $\tilde{m}^{P}_{t}=e^{-r_{p}t}m_{t}^{P}$), we may assume zero interest rates $r_{A}=r_{p}=0$ and to capture the most important feature of Principal-Agent problem, we choose $b(t,a,x)=a,\sigma(t,a,x)=1$. The PA problem constitutes a stochastic control problem; hence, its corresponding Principal’s value function admits characterization via an associated Hamilton-Jacobi-Bellman (HJB) equation. As is standard in the literature (e.g. [sannikov2008continuous]), we first solve the Agent’s problem (2.6) by using the comparison principle of BSDEs and stochastic maximum principle. When solving the Agent’s problem, the Agent’s optimal remaining utility process is characterized as a state process, which we denote as $\{W_{t}\}_{\{0\leq t\leq T\}}$. An auxiliary process $\{Y_{t}\}_{\{0\leq t\leq T\}}$ will also arise when deriving the Agent’s optimal consumption process. These two processes, $\{W_{t}\}$, $\{Y_{t}\}$, together with (2) and (2.4), are the four state processes needed to solve the Principal’s problem.
These states satisfy the following dynamics for $t\leq s\leq T$:
$\begin{cases}W_{s}=&w-\int_{t}^{s}U^{A}(r,\alpha_{r},\beta_{r},I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r}),I_{c}(\alpha_{r},\beta_{r},Y_{r},Z_{r}))dr\\ &+\int_{t}^{s}Z_{r}dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\cr Y_{s}=&y+\int_{t}^{s}\hat{Z}_{r}dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\cr m^{A}_{s}=&\int_{t}^{s}[\alpha_{r}-I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})+\beta_{r}I_{c}(\alpha_{r},\beta_{r},Y_{r},Z_{s})]ds+\int_{t}^{s}\beta_{r}dB^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r}^{A})}_{r},\cr m^{P}_{s}=&\int_{t}^{s}((1-\beta_{r})I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})-D_{r}-\alpha_{r})dr+\int_{t}^{s}(1-\beta_{r})dB_{r}^{I_{a}(\alpha_{r},\beta_{r},Y_{r},Z_{r})},\end{cases}$ (2.9)
@@
$V(t,w,y,0,0)=$ $\sup_{\alpha,\beta,D,Z,\hat{Z}}\mathbb{E}^{I_{a}(\alpha,\beta,Y,Z)}\Big[\int_{t}^{T}U^{P}(s,\alpha_{s},\beta_{s},I_{a}(\alpha_{s},\beta_{s},Y_{s},Z_{s}),D_{s})ds$
$+\Phi^{P}(T,\xi,m^{P}_{T})\Big],$
@@ 2.2 Policy Iteration
In the past, such problems were viewed as very challenging, since they require multiple levels of approximation—locally to solve the supremum and then globally to solve for $V(t,{\bf x})$. Below we develop a deep-learning based setup to address this aspect of PA HJB. The key idea that we rely on is to construct (overparametrized) functional approximator that replaces $V(t,{\bf x})$ with a surrogate $\hat{V}(t,{\bf x})$ that is trained through gradient descent. The approximation is global in $(t,{\bf x})$ (e.g. there is no finite-differencing) and moreover carries through to the derivatives, which are evaluated using algorithmic differentiation. In turn, functional approximation allows to do the same for the optimal control $u(\cdot,\cdot)$, which is similarly approximated via a deep neural network.
The above construction is conceptually dimension-agnostic, and moreover thanks to being overparametrized can be trained exhaustively, breaking down the overall solver into a loop consisting of doing SGD steps over training batches that are refreshed across epochs. Consequently, the method has modest memory requirements and does not rely on any temporal or spatial discretization.
@@
with unknown value function $V$ and unknown optimal feedback control ${\bf u}\equiv u(t,{\bf x})$, with ${\cal L}^{\bf u}V(t,{\bf x})\equiv{\cal L}^{{\bf u}(t,{\bf x})}V(t,{\bf x})$. Here, both $\bf x$} and $\bf u$} are vectors in ${\cal X}\subseteq\mathbb{R}^{d}$ and ${\cal U}\subseteq\mathbb{R}^{a}$ respectively.
@@ 3 Implementation
3.1 Stopping Criteria
Due to the presence of multiple convergence criteria, determining the stopping criterion of our PAAC algorithm is nontrivial. We define the metric
$L_{ctrl}(\theta^{V},\theta^{u};t,{\bf x}):=({\nabla{\cal L}\over\nabla u})^{u(t,{\bf x};\theta^{u})}v(t,{\bf x};\theta^{V})$ (3.16)
to verify whether step 6 in our PAAC algorithm is converging (note that if $\bf u$} is vector, then $L_{ctrl}(\theta^{V},\theta^{u};t,{\bf x})$ is also a vector). To define a stopping criterion, we construct a validation set $\{\bar{t}^{m},\bar{{\bf x}}^{m}\},m=1,\ldots,M_{V}$. Unlike the training batches that vary across steps, the validation set is fixed throughout, to allow an objective tracking of surrogate quality and algorithm convergence and to mitigate over-fitting to a particular training epoch. For the stopping criterion, we seek to control the $L_{\infty}$ norm of the vectors $\bar{L}_{int}$ and $\bar{L}_{ctrl}$ which are counterparts of $L_{int},{L}_{ctrl}$ from (3.16) over the validation set. Thus, we continue the PAAC iterations until we ensure that the maximum pointwise errors of $v_{n},u_{n}$ in the interior are small:
$\|\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot},{\bar{{\bf x}}}^{\cdot})\|_{\infty}\leq\mathrm{Tol_{int}},\qquad\|\bar{L}_{ctrl}(\theta^{V}_{n},\theta^{u}_{n};{\bar{t}}^{\cdot},{\bar{{\bf x}}}^{\cdot})\|_{\infty}\leq\mathrm{Tol_{ctrl}},$
@@ 3.2 Neural Network Architecture
We consider two different architectures for the neural networks. The first one is the DGM (Deep Galerkin Method) architecture used by Sirignano and Spiliopoulos [sirignano2018dgm]. The second one is a residual fully-connected feed-forward neural network.
@@
To evaluate the differential operator ${\cal L}^{{\bf u}(\cdot;\theta^{u}_{n})}$ that defines the loss $L_{int}$ we compute numerical gradients of the current value function approximation $v(\cdot;\theta^{V}_{n})$ using backpropagation. To this end we rely on the native backpropagation (aka algorithmic differentiation) support in modern ML libraries, such as the function tf.gradients in Python’s Tensorflow library. This capability together with the underlying universal approximation property of deep NN’s are the motivating factors for using NN approximators, as it allows to replace “term-for-term” the HJB equation with a corresponding neural network approximation.
@@ 3.3 Training Designs
As a third option, to enhance convergence, we consider adaptive sampling, where new training points are picked based on the current loss $L_{int}(\cdot,\cdot)$, targeting regions where the current loss is highest. This achieves two purposes. First, the use of additional training locations helps to achieve global minimization of the loss metric. Recall that $L_{n}$ is supposed to represent the integrated error over the entire domain; the approximation using a fixed set of $(t^{m},{\bf x}^{m})_{m=1}^{M}$ runs the danger of overfitting at those specific sites. Second, by preferentially picking $(t^{m},{\bf x}^{m})$ we can speed-up convergence by focusing on regions where the present fit is less good.
@@ 3.4 A toy example: The Holmström–Milgrom Model
with exponential utility $\Phi^{A}(x)=-\exp(-\gamma_{A}x)$ for some positive $\gamma_{A}>0$. We denote by $a^{*}(s,\xi)$ the best response in the Agent’s problem. Solving the latter (see proof of Proposition 3.1), we have the terminal payment expressed as
@@
For this example, taking the gradient of the Hamiltonian we have the convergence loss criterion
$L_{ctrl}=\Big[V_{x}^{P}[1-(1+\gamma_{A})Z]+V_{xx}^{P}(Z-1)\Big].$ (3.24)
@@
Figure 2: Estimated optimal control a(t,x){a}(t,x) obtained from the DeepPAAC algorithm for different exponential mixture weights $\lambda\in(0,1)$.
@@
Figure 3: Left: a(t,x;0){a}(t,x;0) for $\lambda=0$. Right: a(t,x;0.5){a}(t,x;0.5) for $\lambda=0.5$
For neural network surrogates of the value function $V(\cdot,\cdot)$ and the control $a(\cdot,\cdot)$, we use our DeepPAAC architecture described in Section 3.2. In the illustration below we take $L=3$ layers, with 32 neurons per layer. The networks $\theta_{0}^{V},\theta_{0}^{u}$ are initialized using Glorot uniform initialization for the weights and zeros initialization for the biases. The stopping rule is $\mathrm{Tol_{int}}=\mathrm{Tol_{ctrl}}=10^{-3}$. For training we use batches of size $M=2000$, polynomial learning rate decay from $10^{-3}$ to $10^{-4}$ with power $0.8$. In the given example we need 6750 steps to satisfy the stopping criterion.
In Figure 2, we plot the numerically obtained feedback control ${a}(t,x;\lambda)$ at a few locations $(t,x)=({1\over 4},{1\over 4}),({1\over 2},{1\over 2}),({3\over 4},{3\over 4})$ as we vary the mixture weight $\lambda$. The parameters in Figure 2 are $\gamma_{A}=0.5$, $\gamma_{P}=1$ and $\tilde{\gamma}_{P}=2$. We observe that ${a}(t,x;\lambda)>a^{lin}(t,x;\lambda)$ while a(t,x;0){a}(t,x;0) and a(t,x;1){a}(t,x;1) match the analytical exact solution. The left panel of Figure 3 shows a(t,x;0){a}(t,x;0) as a function of $t,x$. As expected, the NN surrogate correctly learns that a(t,x;0){a}(t,x;0) is (almost) a constant. The right panel of Figure 3 repeats this for a(t,x;0.5){a}(t,x;0.5). We observe that when $\lambda=0.5$, ${a}(t,x;\lambda)$ increases in $t$ and decreases in $x$.